Differentiate i^i...

Grade:12

2 Answers

Akash Kumar Dutta
11 years ago

i^i = e^-(pi/2)   [ by eularian form]
hence d/dx e^-(pi)/2= 0(ANS)

Abhishekh kumar sharma
10 years ago

in eulerian form i=e^∏\2.Φ=a let

then put i^i=a^a

then diffrentiate

easy